OpenStudy (anonymous):

Help I Need Help On This Problem . . . Which of the following best describes the solution to the system of equations below? 7x + 6y=6 7x + 6y=14 A. The system of equations has exactly one solution where x= 0 and y=1 . B. The system of equations has exactly one solution where x=6 and y=14 . C. The system of equations has no solution. D. The system of equations has infinitely many solutions.

OpenStudy (joshawad):

7x+ 6y=6 7x x 6y 42xy=6 7xy= 1

OpenStudy (anonymous):

D is the answer 7x + 6y = 6 -7x -6y = -14 _______________ 0 + 0 = -8 (or 0 = 8) not possible

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Is Correct